# Linear-quadratic optimal control for abstract differential-algebraic equations

### v1

@@ Abstract
In this paper, we extend classical approach to linear quadratic (LQ) optimal control via Popov operators to abstract linear differential-algebraic equations in Hilbert spaces. To ensure existence of solutions, we assume that the underlying differential-algebraic equation has index one in the pseudo-resolvent sense. This leads to the existence of a degenerate semigroup that can be used to define a Popov operator for our system. It is shown that under a suitable coercivity assumption for the Popov operator the optimal costs can be described by a bounded Riccati operator and that the optimal control input is of feedback form. Furthermore, we characterize exponential stability of abstract differential-algebraic equations which is required to solve the infinite horizon LQ problem.
Keywords: optimal control, differential-algebraic equations, infinite dimensional systems, linear quadratic regulators, stability
@@ 1 Introduction
The specific challenge arises from the fact that, in numerous practically motivated instances, as observed in [MZ23, Rei07, Rei06, RT05], the operator $E$ possesses non-trivial kernel and co-kernel. Examples of such scenarios include abstract Cauchy problems incorporating additional closure relations, physical problems characterized by vanishing material parameters or densities, or systems of coupled partial differential equations (PDEs) like heat-wave couplings formulated in an abstract differential-algebraic form. Subject to the dynamics (1), our objective is to minimize the quadratic cost on some time horizon $t_{f}>0$ given by
@@
where $\mathcal{Q}\in L^{\infty}([0,t_{f}],L(Y))$, $\mathcal{R}\in L^{\infty}([0,t_{f}],L(U))$, $\mathcal{N}\in L^{\infty}([0,t_{f}],L(Y,U))$ (that is, $\mathcal{Q}$, $\mathcal{R}$ and $\mathcal{N}$ are bounded functions with values in the space of linear operators), and additionally, $Q$ and $R$ are pointwise self-adjoint. This minimization is performed over all solutions of (1) with $u\in L^{2}([0,t_{f}],U)$ and $y\in L^{2}([0,t_{f}],Y)$. As a main assumption, we will use that the abstract differential-algebraic system has index at most one and is stable in the pseudo-resolvent sense, meaning that the following holds
@@
The incorporation of this index condition, alongside an additional stability assumption, enables us to relate our setup to the optimal control framework for standard infinite-dimensional systems presented in [WW97].
We will further assume that the transfer function $G(s)=C(sE-A)^{-1}B$ is bounded in the complex right half-plane, that is
@@
Our findings are inspired by previous results for finite-dimensional differential-algebraic equations [RV19], see also [Meh91] for an overview in the ODE case.
@@ 2 Preliminaries
In this section, we recall some results from [GR23] on solvability of abstract differential-algebraic equations of the form
@@ 2.1 Pseudo-Resolvents
$(\mu-\lambda)(A-\mu E)^{-1}E(A-\lambda E)^{-1}$ $=(A-\lambda E)^{-1}-(A-\mu E)^{-1}$ (12)
@@
Multiplying (12) by $E$ from both the left and the right implies that
@@
are what are known as left and right pseudo-resolvents, see [GR23]. The openness of $\Omega:=\rho(E,A)$, combined with the identity above, implies that both $R_{l}$ and $R_{r}$ are holomorphic operator-valued functions.
@@
$\|R(\lambda)\|\leq\frac{M}{\lambda-\omega},\quad\lambda\in\Omega\cap\mathbb{R}\text{ with }\lambda>\omega.$ (15)
then by a generalization of the Hille-Yosida theorem it generates a so-called degenerate semigroup, see [Are01, FY99, GR23]. This concept will be treated in the next section.
We first show that pseudo-resolvents with property (15) also admit a decomposition of $X$, see e.g. [Kat59].
@@
Let $X$ be a Hilbert space and $R:\Omega\rightarrow L(X)$ be a pseudo-resolvent which satisfies (15) then for all $\lambda\in\Omega$, $X$ decomposes into the direct sum
$X=X_{K}\dotplus X_{R}=\ker R(\lambda)\dotplus\overline{{\rm ran}\,R(\lambda)}$ (16)
@@
It was shown in [JM22, Theorem 3.6], that the growth bound follows from dissipativity conditions on the coefficients $E$ and $A$, see also [GR23, Section 7] for a possible extension to a Banach space setting and a more general dissipativity concept.
@@
If $\lambda\in\rho(E,A)$ for some $\lambda>0$ and
@@
$\operatorname{Re}\langle A^{*}x,E^{*}x\rangle_{X}$ $\leq 0,x\in D(A^{*}),$
@@ 2.2 Degenerate semigroups
Since $X$ is a Hilbert space, it is reflexive and this implies, see [Are01, Corollary 2.2] that the following strong limit exists $T(0):=\lim\limits_{t\downarrow 0}T(t)$. Degenerate semigroups with this property are called strongly continuous. In this case $T(0)$ is a bounded projection onto $X_{R}:={\rm ran}\,T(0)$. The restriction of $T$ to $X_{R}$, i.e., $(T(t)|_{X_{R}})_{t\geq 0}$, is moreover a strongly continuous semigroup on $X_{R}$. Further, we have that $\ker T(t)=\ker T(0)$ for all $t\geq 0$. In particular,
@@
${\rm graph}\,A_{R}=\{(R(\lambda)x,x+\lambda R(\lambda)x),x\in\overline{{\rm ran}\,R(\lambda)}\}\subseteq X_{R}\times X_{R}.$ (17)
@@
Hence according to [ABHN11, Corollary 3.7.12] it generates a $C_{0}$-semigroup $(T_{R}(t))_{t\geq 0}$ on $X_{R}$. Hence, we obtain a degenerate semigroup for $x=x_{0}+x_{1}\in X_{K}\dotplus X_{R}$
@@ 2.3 Exponential Stability of semigroups
We recall the following characterization of exponential stability of semigroups, see e.g. [JZ12, Theorem 8.1.3, Lemma 8.1.2, Theorem 8.1.4].
@@
Let $A$ be the infinitesimal generator of the $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on the Hilbert space $X$. Then the following is equivalent.
@@
    $\langle AQ,Qx\rangle+\langle Qx,Ax\rangle=-\langle x,x\rangle,\quad\text{for all $x\in D(A)$.}$
@@ 2.4 Mild Solutions
Motivated by the study of linear quadratic optimal control problems, we consider here mild solutions of abstract differential-algebraic equations. Classical and weak solutions are discussed in further detail in [GR23]. We call $x\in L^{2}([0,t_{f}],X)$ is a mild solution of (11) on $[0,t_{f})$, $t_{f}\in[0,\infty]$, if $Ex:[0,t_{f})\to Z$ is continuous, and
@@
$(A-\mu E)^{-1}A=(A-\mu E)^{-1}(A-\mu E+\mu E)=I-\mu(A-\mu E)^{-1}E,$
we obtain, for $\hat{f}=(A-\lambda E)^{-1}f$.
$\tfrac{{\rm d}}{{\rm d}t}R_{r}(\mu)x(t)=(I-\mu R_{r}(\mu))x(t)+\hat{f}(t),\quad R_{r}(\mu)x(0)=R_{r}(\mu)x_{0}.$
In particular, if $0\in\rho(E,A)$, we have, for $\hat{f}=A^{-1}f$.
@@
To describe the mild solutions explicitly, we decompose $\hat{f}=\hat{f}_{R}+\hat{f}_{K}\in L^{2}([0,t_{f}],X)$ according to the direct sum decomposition $X=X_{K}\dotplus X_{r}=\ker R(0)\dotplus\overline{{\rm ran}\,R(0)}$. Since $X_{K}$ and $X_{R}$ are both closed, there exists a bounded projector $P\in L(X)$ with ${\rm ran}\,P=X_{R}$ and $\ker P={\rm ran}\,(I-P)=X_{K}$.
@@
$\hat{f}=\hat{f}_{R}+\hat{f}_{K}=P\hat{f}+(I-P)\hat{f}.$
The mild solutions of the pseudo-resolvent equation (21) have been studied in [GR23, Lemma 8.1, Proposition 8.3]. Below, we state the result only for the pseudo-resolvent $R_{r}(\lambda)=(A-\lambda E)^{-1}E$ and in the case that $\mu=0\in\rho(A)$.
We first present some facts from [TW09] on unbounded control operators: For generator $D(A_{R})\subset X_{R}\to X_{R}$ of a strongly continuous semigroup $(T_{R}(t))_{t\geq 0}$, we can associate the space $X_{R,-1}$, which is obtained by the completion of $X$ with respect the norm
@@
Further, $A_{R}$ extends to an operator $A_{R,-1}$, which generated a semigroup $(T_{R,-1}(t))_{t\geq 0}$ on $X_{R,-1}$. The latter is also an extension of $(T_{R}(t))_{t\geq 0}$, and it it referred to as extrapolated semigroup. If $f_{R}:[0,t_{f}]\to X_{R,-1}$ is weakly differentiable, then the mild solution of $\tfrac{{\rm d}}{{\rm d}t}x_{R}(t)=A_{R}x_{R}(t)+f_{R}(t)$, $x_{R}(0)=x_{R,0}\in X_{R}$ is given by $x:[0,t_{f}]\to X_{R}$ with
@@
$T_{-1}(t):\quad X_{R,-1}\dotplus X_{K}\to X_{R,-1}\dotplus X_{K},\quad x_{0,R}+x_{0,K}\mapsto T_{R,-1}(t)x_{0,R}.$
@@
Given the abstract differential-algebraic equation (11) such that $(E,A)$ has pseudo resolvent index one and some $\omega\in\mathbb{R}$ and let $\omega<0$ and let $T$ be the degenerated semigroup associated to $R(\lambda)=(A-\lambda E)^{-1}E$, and let $T_{-1}$ be the corresponding extrapolated semigroup. Let $A_{R}:D(A_{R})\subset X_{R}\to X_{R}$ be the operator as in (17). Further, decompose $\hat{f}=\hat{f}_{R}+\hat{f}_{K}\in L^{2}([0,t_{f}],X)$ with $\hat{f}_{R}\in L^{2}([0,t_{f}],\overline{{\rm ran}\,R(0)})$ and $\hat{f}_{K}\in L^{2}([0,t_{f}],\ker R(0))$. Then for $x_{0}\in\overline{{\rm ran}\,R(0)}$ the unique mild solution of (11) is given by
$x(t)=T_{R}(t)x_{0}+\int_{0}^{t}T_{R,-1}(t-\tau){A_{R}}\hat{f}_{R}(s)\,{\rm d}\tau-\hat{f}_{K}(t),$ (22)
@@
First, we show that the mild solution of the pseudo-resolvent equation (21) coincide with the mild solutions of (11). Let $A^{-1}Ex:[0,t_{f})\rightarrow X$ be continuous and satisfying for all $t\geq 0$
$A^{-1}Ex(t)-A^{-1}Ex(0)=\int_{0}^{t}x(\tau)\,{\rm d}\tau+\int_{0}^{t}\hat{f}(\tau)\,{\rm d}\tau=\int_{0}^{t}x(\tau)\,{\rm d}\tau+A^{-1}\int_{0}^{t}f(\tau)\,{\rm d}\tau.$ (23)
Since ${\rm ran}\,A^{-1}=D(A)$, we conclude $\int_{0}^{t}x(\tau)\,{\rm d}\tau\in D(A)$. Therefore, applying $A$ to (23) leads to (20) and therefore $x$ is a mild solution of (11). Conversely, if $x$ is a mild solution of (11), then $Ex$ is continuous and therefore $A^{-1}Ex$ is continuous and multiplying with $A^{-1}$ from the left in (20) implies (23).
We proceed by showing that $x$ given by (22) is the unique mild solution of the pseudo-resolvent equation (21). To this end, we construct a mild solution $x_{R}$ on $X_{R}$ and $x_{K}$ on $X_{K}$. First, observe that $x_{K}(t):=-\hat{f}_{K}\in L^{2}([0,t_{f}],X_{K})$ satisfies (23) pointwise for all $t\in[0,t_{f}]$. The mild solutions $x_{R}$ of (23) for $x_{R}(0)=x_{R}^{0}\in X_{R}$
$A^{-1}Ex_{R}(t)-A^{-1}Ex_{R}(0)=\int_{0}^{t}x_{R}(\tau)\,{\rm d}\tau+\int_{0}^{t}PA^{-1}f(\tau)\,{\rm d}\tau.$ (24)
This can be rewritten using the operator $A_{R}$ from (17) which is invertible and fulfills $A_{R}^{-1}=(A^{-1}E)|_{X_{R}}$. Therefore, (24) is equivalent to
@@
Since $A_{R}$ generates by assumption the $C_{0}$-semigroup $\{T_{R}(t)\}_{t\geq 0}$ it follows that the unique mild solution $x_{R}$ of (25) is given by

### v2

@@ Abstract
In this paper, we extend a classical approach to linear quadratic (LQ) optimal control via Popov operators to abstract linear differential-algebraic equations (ADAEs) in Hilbert spaces. To ensure existence of solutions, we assume that the underlying differential-algebraic equation has index one in the pseudo-resolvent sense. This leads to the existence of a degenerate semigroup that can be used to define a Popov operator for our system. It is shown that under a suitable coercivity assumption for the Popov operator the optimal costs can be described by a bounded Riccati operator and that the optimal control input is of feedback form. Furthermore, we characterize exponential stability of ADAEs which is required to solve the infinite horizon LQ problem.
Keywords: optimal control, differential-algebraic equations, infinite dimensional systems, linear quadratic regulator, stability
@@ 1 Introduction
The specific challenge arises from the fact that, in numerous practically motivated instances, as observed in [11, 15, 14, 16], the operator $E$ possesses non-trivial kernel and co-kernel. Examples of such scenarios include abstract Cauchy problems incorporating additional closure relations, physical problems characterized by vanishing material parameters or densities, or systems of coupled partial differential equations (PDEs) like heat-wave couplings formulated in an abstract differential-algebraic form. Subject to the dynamics (1), our objective is to minimize the quadratic cost on some time horizon $t_{f}>0$ given by
@@
where $\mathcal{Q}\in L^{\infty}([0,t_{f}],L(Y))$, $\mathcal{R}\in L^{\infty}([0,t_{f}],L(U))$, $\mathcal{N}\in L^{\infty}([0,t_{f}],L(Y,U))$ (that is, $\mathcal{Q}$, $\mathcal{R}$ and $\mathcal{N}$ are bounded functions with values in the space of linear operators), and additionally, $Q$ and $R$ are pointwise self-adjoint. This minimization is performed over all solutions of (1) with $u\in L^{2}([0,t_{f}],U)$ and $y\in L^{2}([0,t_{f}],Y)$.
As a main assumption, we use that the abstract differential-algebraic system has index at most one and is stable in the pseudo-resolvent sense, meaning that the following holds
@@
The incorporation of this index condition, alongside an additional stability assumption, enables us to relate our setup to the optimal control framework for standard infinite-dimensional systems presented in [20].
Further, we assume that the transfer function $G(s)=C(sE-A)^{-1}B$ is bounded in the complex right half-plane, that is
@@
Our findings are inspired by previous results for finite-dimensional differential-algebraic equations [17], see also [12] for an overview in the ODE case. Furthermore, we use the solution theory that is developed for inhomogeneous ADAEs of finite pseudo-resolvent index [7] to define Popov operators for ADAEs. These operators can be used so solve the LQ problem on finite-time and infinite time-horizons. Recently, a different approach to LQ optimal control of ADAEs with radiality index was presented in [1] using the solutions of differential Riccati equations, see also [5, 7] for a comparison of the different index concepts.
In this note, we consider also the infinite horizon LQ problem, and here it is necessary to have exponential stability of the underlying degenerate semi-groups. Therefore, we also provide Lyapunov-like characterizations of exponential stability in terms of the pseudo-resolvents and in terms of the coefficients of the ADAE.
@@ 2 Preliminaries
In this section, we recall some results from [7] on solvability of abstract differential-algebraic equations of the form
@@ 2.1 Pseudo-Resolvents
$\,\,\,\,\,\,\,\,(\mu-\lambda)(A-\mu E)^{-1}E(A-\lambda E)^{-1}$
$=(A-\lambda E)^{-1}-(A-\mu E)^{-1}$ (6)
@@
Multiplying (2.1) by $E$ from both the left and the right implies that
@@
are what are known as left and right pseudo-resolvents, see [7]. The openness of $\Omega:=\rho(E,A)$, combined with the identity above, implies that both $R_{l}$ and $R_{r}$ are holomorphic operator-valued functions.
@@
$\|R(\lambda)\|\leq\frac{M}{\lambda-\omega},\quad\lambda\in\Omega\cap\mathbb{R}\text{ with }\lambda>\omega,$ (9)
then by a generalization of the Hille-Yosida theorem it generates a so-called degenerate semigroup, see [2, 7]. This concept will be treated in the next section.
We first show that pseudo-resolvents with property (9) also admit a decomposition of $X$, see e.g. [10].
@@
Let $X$ be a Hilbert space and $R:\Omega\rightarrow L(X)$ be a pseudo-resolvent which satisfies (9). Then for all $\lambda\in\Omega$, $X$ decomposes into the direct sum
$X=X_{K}\dotplus X_{R}=\ker R(\lambda)\dotplus\overline{{\rm ran}\,R(\lambda)},$ (10)
@@
It is shown in [8, Theorem 3.6], that the growth bound (9) follows from dissipativity conditions on the coefficients $E$ and $A$, see also [7, Section 7] for a possible extension to a Banach space setting and a more general dissipativity concept.
@@
Let $\lambda\in\rho(E,A)$ for some $\lambda>0$ and assume that
@@
$\operatorname{Re}\langle A^{*}x,E^{*}x\rangle_{X}$ $\leq 0,x\in D(A^{*}).$
@@ 2.2 Degenerate semigroups
Since $X$ is a Hilbert space, it is reflexive and this implies, see [2, Corollary 2.2] that the following strong limit exists $T(0):=\lim\limits_{t\downarrow 0}T(t)$. Degenerate semigroups with this property are called strongly continuous. In this case $T(0)$ is a bounded projection onto $X_{R}:={\rm ran}\,T(0)$. The restriction of $T$ to $X_{R}$, i.e., $(T(t)|_{X_{R}})_{t\geq 0}$, is moreover a strongly continuous semigroup on $X_{R}$. Further, we have that $\ker T(t)=\ker T(0)$ for all $t\geq 0$. In particular,
@@
${\rm graph}\,A_{R}\\ =\{(R(\lambda)x,x+\lambda R(\lambda)x),x\in\overline{{\rm ran}\,R(\lambda)}\}\subseteq X_{R}\times X_{R}.$ (11)
@@
Hence, according to [3, Corollary 3.7.12] it generates a $C_{0}$-semigroup $(T_{R}(t))_{t\geq 0}$ on $X_{R}$. Hence, we obtain a degenerate semigroup for $x=x_{0}+x_{1}\in X_{K}\dotplus X_{R}$
@@ 2.3 Exponential Stability of semigroups
We recall the following characterization of exponential stability of semigroups, see e.g. [9, Theorem 8.1.3, Lemma 8.1.2, Theorem 8.1.4].
@@
Let $A$ be the infinitesimal generator of the $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on the Hilbert space $X$. Then the following is equivalent
@@
    $\langle Ax,Qx\rangle+\langle Qx,Ax\rangle=-\langle x,x\rangle,\quad\text{for all $x\in D(A)$.}$
@@ 2.4 Mild Solutions
Motivated by the study of linear quadratic optimal control problems, we consider here mild solutions of abstract differential-algebraic equations. Classical and weak solutions are discussed in further detail in [7]. We call $x\in L^{2}([0,t_{f}],X)$mild solution of (5) on $[0,t_{f}]$, $t_{f}>0$, if $Ex:[0,t_{f}]\to Z$ is continuous, and
@@
$(A-\mu E)^{-1}A=(A-\mu E)^{-1}(A-\mu E+\mu E)\\ =I-\mu(A-\mu E)^{-1}E,$
we obtain, for $\hat{f}=(A-\lambda E)^{-1}f$
$\tfrac{{\rm d}}{{\rm d}t}R_{r}(\mu)x(t)$ $=(I-\mu R_{r}(\mu))x(t)+\hat{f}(t),$
$R_{r}(\mu)x(0)$ $=R_{r}(\mu)x_{0}.$
In particular, if $0\in\rho(E,A)$, we have, for $\hat{f}=A^{-1}f$ that
@@
To describe the mild solutions explicitly, we decompose $\hat{f}=\hat{f}_{R}+\hat{f}_{K}\in L^{2}([0,t_{f}],X)$ according to the direct sum decomposition $X=X_{K}\dotplus X_{R}=\ker R(0)\dotplus\overline{{\rm ran}\,R(0)}$. Since $X_{K}$ and $X_{R}$ are both closed, there exists a bounded projector $P\in L(X)$ that satisfies
${\rm ran}\,P=X_{R},\quad\ker P={\rm ran}\,(I-P)=X_{K}.$ (16)
@@
$A^{-1}f=\hat{f}=\hat{f}_{R}+\hat{f}_{K}=P\hat{f}+(I-P)\hat{f}.$ (17)
The mild solutions of the pseudo-resolvent equation (15) have been studied in [7, Lemma 8.1, Proposition 8.3]. Below, we state the result only for the pseudo-resolvent $R_{r}(\lambda)=(A-\lambda E)^{-1}E$ and in the case that $\mu=0\in\rho(A)$.
We first present some facts from [18] on unbounded control operators: For generator $D(A_{R})\subset X_{R}\to X_{R}$ of a strongly continuous semigroup $(T_{R}(t))_{t\geq 0}$, we can associate the space $X_{R,-1}$, which is obtained by the completion of $X$ with respect to the norm
@@
Further, $A_{R}$ extends to an operator $A_{R,-1}$, which generates a semigroup $(T_{R,-1}(t))_{t\geq 0}$ on $X_{R,-1}$. The latter is also an extension of $(T_{R}(t))_{t\geq 0}$, and it it referred to as extrapolated semigroup. If $f_{R}:[0,t_{f}]\to X_{R,-1}$ is weakly differentiable, then the mild solution of $\tfrac{{\rm d}}{{\rm d}t}x_{R}(t)=A_{R}x_{R}(t)+f_{R}(t)$, $x_{R}(0)=x_{R,0}\in X_{R}$ is given by $x:[0,t_{f}]\to X_{R}$ with
@@
$T_{-1}(t):\quad X_{R,-1}\dotplus X_{K}$ $\to X_{R,-1}\dotplus X_{K},$
$x_{0,R}+x_{0,K}$ $\mapsto T_{R,-1}(t)x_{0,R}.$
@@
Given the abstract differential-algebraic equation (5) such that $R_{r}$ given by (8) satisfies (3) for some $\omega<0$ and let $T$ be the degenerated semigroup associated to $R(\lambda)=(A-\lambda E)^{-1}E$, and let $T_{-1}$ be the corresponding extrapolated semigroup. Let $A_{R}:D(A_{R})\subset X_{R}\to X_{R}$ be the operator as in (11). Further, we decompose $\hat{f}=\hat{f}_{R}+\hat{f}_{K}\in L^{2}([0,t_{f}],X)$ with $\hat{f}_{R}\in L^{2}([0,t_{f}],\overline{{\rm ran}\,R(0)})$ and $\hat{f}_{K}\in L^{2}([0,t_{f}],\ker R(0))$ as in (17). Then for $x_{0}\in\overline{{\rm ran}\,R(0)}$ the unique mild solution of (5) is given by
$x(t)=T_{R}(t)x_{0}+\int_{0}^{t}T_{R,-1}(t-\tau){A_{R}}\hat{f}_{R}(s)\,{\rm d}\tau-\hat{f}_{K}(t).$ (18)
@@
First, we show that $x$ given by (18) is the unique mild solution of the pseudo-resolvent equation (15). To this end, we construct a mild solution $x_{R}$ on $X_{R}$ and $x_{K}$ on $X_{K}$. First, observe that $x_{K}(t):=-\hat{f}_{K}\in L^{2}([0,t_{f}],X_{K})$ satisfies (21) pointwise for all $t\in[0,t_{f}]$. The mild solutions $x_{R}$ of (21) for $x_{R}(0)=x_{R}^{0}\in X_{R}$
$A^{-1}Ex_{R}(t)-A^{-1}Ex_{R}(0)=\int_{0}^{t}(x_{R}(\tau)+PA^{-1}f(\tau))\,{\rm d}\tau,$ (19)
where $P$ is the projection onto $X_{R}$ that is defined in (16). This can be rewritten using the operator $A_{R}$ from (11), which is invertible and fulfills $A_{R}^{-1}=(A^{-1}E)|_{X_{R}}$. Therefore, (19) is equivalent to
@@
By assumption $A_{R}$ generates the $C_{0}$-semigroup $\{T_{R}(t)\}_{t\geq 0}$, and hence, the unique mild solution $x_{R}$ of (20) is given by
@@
Next, we show that the mild solution of the pseudo-resolvent equation (15) coincide with the mild solutions of (5). Let $A^{-1}Ex:[0,t_{f})\rightarrow X$ be continuous and satisfying for all $t\geq 0$
$A^{-1}Ex(t)-A^{-1}Ex(0)$ $=\int_{0}^{t}x(\tau)\,{\rm d}\tau+\int_{0}^{t}\hat{f}(\tau)\,{\rm d}\tau$ (21)
$=\int_{0}^{t}x(\tau)\,{\rm d}\tau+A^{-1}\int_{0}^{t}f(\tau)\,{\rm d}\tau.$
Since ${\rm ran}\,A^{-1}=D(A)$, we conclude $\int_{0}^{t}x(\tau)\,{\rm d}\tau\in D(A)$. Therefore, applying $A$ to (21) leads to (14) and therefore $x$ is a mild solution of (5). Conversely, if $x$ is a mild solution of (5), then $Ex$ is continuous and therefore $A^{-1}Ex$ is continuous and multiplying with $A^{-1}$ from the left in (14) implies (21). ∎∎